Check List for Particle : d̄

April 21, 2009

Quantity  Us  PDG  EvtGen  Quantity  Us  PDG  EvtGen
PDG  -1  Mass  5

Code

Width 0 Gen dbar

Name

Decay  0 m.

Length

Latex Name d̄  Mass Gen.  Width Gen.  Type quark
Type  constituent  Con. Mass 325  Rating  ∗ ∗ ∗ ∗ ∗  Variable  fixed

Ratio

Stable  stable  Spin  1/2  Charge  1/3  Colour  -3
Lower Cut  0  Upper cut  0  PDG link  q123

• Is the description present and correct? The down quark is the isospin=−1/2 first generation quark. The physical mass is at the centre of the range quoted in [?] and is only used as a coupling while the constituent mass is used kinematically.

• References?

Branching Rating Outgoing Description  Decayer EvtGen
Ratio  Particles

Table 1: The decay modes of the d̄.

Total branching ratio is 0.